OpenStudy (anonymous):

Solve the equation Quantity x minus 5 all over 4 = 3 for x. Please show your work

jhonyy9 (jhonyy9):

(x-5)/4 =3/x is cirrect ?

OpenStudy (anonymous):

3-5/4 ?

jhonyy9 (jhonyy9):

so than can you please cross multiplie ?

jhonyy9 (jhonyy9):

ohhh yes (x-5)/4 =3 yes ?

OpenStudy (anonymous):

(x - 5)/4 x - 5 = 12 x = 17

jhonyy9 (jhonyy9):

yes is right